# ATLAS Photometry of Interstellar Object 3I/ATLAS

### v1

@@ Abstract
We present calibrated ATLAS photometry of the interstellar comet 3I/ATLAS (C/2025 N1) from 28 March through 29 Aug 2025, obtained with the five-site, robotic ATLAS network in the $c$ (420–650 nm), $o$ (560–820 nm), and Teide $w$ (420-720 nm) bands. Stacked difference images yield reliable light curves measured in four fixed apertures that capture the evolving coma. We observe 3I/ATLAS transitioning in color from red $(c-o)\approx 0.7$ before MJD 60860 to near-solar $(c-o)\approx 0.4$ after MJD 60870, coincident with the appearance of a prominent anti-solar tail. The absolute magnitude curve $H(t)$ shows a slope break near MJD 60890 at $r\sim 3.3$ au from 0.035-0.035 to 0.014-0.014 mag/day, or in terms of coma cross section as a function of heliocentric distance, $r^{-3.9}$ to $r^{-1.2}$. We interpret the evolution as a shift from dust lifted from a reddened surface to production of small, optically bright icy grains and changing coma optical depth. We release the aperture photometry with geometry and uncertainties to enable cross-instrument synthesis of 3I/ATLAS activity and color evolution.
@@ 1 Introduction
Seligman et al. (2025) presents a calculation that estimates the ATLAS discovery cross section multiplied by typical interstellar object velocity as $600$ au3/yr for objects of the $\sim$10 km effective size of 3I/ATLAS including its coma, so the detection of two such objects in 6 years by ATLAS implies a local density of $\sim$$3\times 10^{-4}$ au-3. The volume inside of Jupiter probably contains a new $\sim$10 km (effective) scale interstellar object every 5 years.
@@
The subject of this paper is more mundane than these population questions, however. Since ATLAS observes all the sky all the time we have likely detections of 3I/ATLAS as early as late April 2025, and continuously ever since. During this span of time the 3I/ATLAS coma has changed dramatically from a red “rubble fountain” pointing towards the sun to a fully developed, white comet tail pointing away. The ATLAS instruments do not have the resolution to isolate the comet nucleus (even HST and JWST cannot do that), so we present the ATLAS photometry through 30 Aug 2025 in a set of 4 concentric apertures that capture more and more coma brightness.
@@ 2 Observations and Data
We normally collect 4 images each night of each field because linkage of moving objects is much simpler if 4 detections exist within a short period of time, although there are certainly methods to link over longer time intervals (Tonry, 2023; Heinze et al., 2022; Holman et al., 2018). Images of objects that vary more slowly than the $\sim$40 min spent collecting the data can be rebinned onto a common coordinate system (moving for a moving object) and stacked for improved signal to noise. The difference imaging stage is particularly important in crowded fields such as the galactic plane where 3I/ATLAS was discovered, but it requires great care in astrometric solutions, $\sim$0.02 pixel or better. Similarly, the stacking of the earliest images from ATLAS had to wait for the extrapolation of the 3I/ATLAS orbit to attain sub-arcsecond accuracy.
Figures 1 and 2 show the appearance of 3I/ATLAS in the stacked images.
@@
Figure 1: The discovery $o$ band image of 3I/ATLAS from MJD 60957 is shown on the left centered on 3I/ATLAS, and the four difference images from that night are stacked on the right. The 9.3′ image on the left has north down with the native 1.86″pixels; the 10′ image on the right has north up with 2″ pixels; black indicates surface brightness less than 21 AB arcsec-2.
@@
Figure 2: These images illustrate the evolution of 3I/ATLAS. From left to right are the $c$ band stack from MJD 60818–68026 (28 May) the $o$ band stack from 60817–60828 (26 May), the epoch at 60884 (28 Jul) just before the $H$ band slope change described below, and the epoch at 60905 (17 Aug) when the eastward, anti-solar tail becomes prominent. The fields are 100″ with 2″ pixels, north up and east left, black indicates surface brightness less than 23.2 AB ″-2.
@@
Unfortunately most of the photometry reported in the literature does not provide enough information for intercomparison. In this paper we use $c$ and $o$ bandpasses whose detailed profiles are published, we specify the reference to which our photometry zeropoints are tied, the apertures for the photometry are exactly defined (although seeing obviously plays a role in smaller apertures), we illustrate how our difference images credibly remove stellar contamination, and extremely importantly, our background level is derived from a constant plus $\rho^{-3}$ fit to median ring surface brightnesses out to a radius of $\rho\sim 50$″.
@@
Table LABEL:tab:phot gives the cumulative magnitude measured for 3I/ATLAS in four apertures drawn from a grid of 2″ pixels. “StackID” codes site, instrument, and integer MJD night number, where a hyphen indicates the MJD range if more than $\sim$1 hour, “MJD” is the actual mean time of the exposure (1 July 2025 is MJD 60857), “N” is the number of exposures contributing to the stack, “m2” is the central 2″ pixel, “m6” is the 9 pixel $\times6$″ square aperture, “m10” adds to “m6” the nearest 12 pixels to make a rounded aperture of diameter $\sim$10″, “m14” adds to “m10” the nearest 16 pixels to make a rounded aperture of diameter $\sim$14″, “dm6” is the uncertainty of the “m6” magnitude (other uncertainties can be scaled by square root of area), “r” is the heliocentric distance [au], “delta” is the Earth distance [au], “sto” is the Sun-Target-Observer angle [deg], and $\Phi$ is the standard $G=0.15$ phase function evaluated at sto. The asterisk on 05r60868 and 05r60883 indicates potentially unreliable photometry, as described below. The typical point spread function for these observations is about 4″ full width half maximum (the $o$ and $c$ images have 1.86″/pixel and the $w$ band 1.25″/pixel) so the central pixel is significantly diminished by the seeing relative to infinite resolution. These data are available in at http://astroportal.ifa.hawaii.edu/atlas/3I-photometry.

### v3

@@ Abstract
We present calibrated ATLAS photometry of the interstellar comet 3I/ATLAS (C/2025 N1) from 28 March through 29 Aug 2025, obtained with the five-site, robotic ATLAS network in the $c$ (420-650 nm), $o$ (560-820 nm), and Teide $w$ (420-720 nm) bands. Stacked difference images yield reliable light curves measured in four fixed apertures that capture the evolving coma. We observe 3I/ATLAS transitioning in color from red $(c-o)\approx 0.7$ before MJD 60860 to near-solar $(c-o)\approx 0.3$ after MJD 60870, coincident with the appearance of a prominent anti-solar tail. The absolute magnitude curve $H(t)$ shows a slope break near MJD 60890 at $r\sim 3.3$ au from 0.035-0.035 to 0.012-0.012 mag/day, or in terms of coma cross section as a function of heliocentric distance, $r^{-3.9}$ to $r^{-1.1}$. We release the aperture photometry with geometry and uncertainties to enable cross-instrument synthesis of 3I/ATLAS activity and color evolution.
@@ 1 Introduction
We presented a calculation (Seligman et al., 2025) that estimates the ATLAS discovery cross section multiplied by typical interstellar object velocity as $600$ au3/yr for objects of the $\sim$10 km effective size of 3I/ATLAS including its coma, so the detection of two such objects in 6 years by ATLAS implies a local density of $\sim$$3\times 10^{-4}$ au-3. The volume inside of Jupiter probably contains a new $\sim$10 km (effective) scale interstellar object every 5 years.
@@
The subject of this paper is more mundane than these population questions, however. Since ATLAS observes all the sky all the time we have likely detections of 3I/ATLAS as early as April 2025, and continuously ever since. During this span of time the 3I/ATLAS coma has changed dramatically from a red “rubble fountain” pointing towards the sun to a fully developed, white comet tail pointing away. The ATLAS instruments do not have the resolution to isolate the comet nucleus (even HST and JWST cannot do that), so we present the ATLAS photometry through 30 Aug 2025 in a set of 4 concentric apertures that capture more and more coma brightness.
@@ 2 Observations and Data
We normally collect 4 images each night of each field because linkage of moving objects is much simpler if 4 detections exist within a short period of time, although there are certainly methods to link over longer time intervals (Tonry, 2023; Heinze et al., 2022; Holman et al., 2018). Images of objects that vary more slowly than the $\sim$40 min spent collecting the data can be rebinned onto a common coordinate system (moving for a moving object) and stacked for improved signal to noise. The difference imaging stage is particularly important in crowded fields such as the galactic plane where 3I/ATLAS was discovered, but it requires great care in astrometric solutions, $\sim$0.02 pixel or better. Similarly, the stacking of the earliest images from ATLAS had to wait for the extrapolation of the 3I/ATLAS orbit to attain sub-arcsecond accuracy. Figures 1 and 2 show the appearance of 3I/ATLAS in the stacked images.
@@
Figure 1: The discovery $o$ band image of 3I/ATLAS from MJD 60957 is shown on the left centered on 3I/ATLAS, and the four difference images from that night are stacked on the right. The 9.3′ image on the left has north up, east left with the native 1.86″pixels; the 10′ image on the right has 2″ rebinned pixels; black indicates surface brightness brighter than 21 AB arcsec-2.
@@
Figure 2: These images illustrate the evolution of 3I/ATLAS. From left to right are the $c$ band stack from MJD 60818–60826 (28 May) the $o$ band stack from 60817–60828 (26 May), the epoch at 60884 (27 Jul) just before the $H$ band slope change described below, and the epoch at 60905 (17 Aug) when the eastward, anti-solar tail becomes prominent. The fields are 100″ with 2″ pixels, north up and east left, black indicates surface brightness brighter than 23.2 AB arcsec-2.
@@
Unfortunately most of the photometry reported in the literature does not provide enough information for intercomparison. In this paper we use $c$ and $o$ bandpasses whose detailed profiles are published, we specify the reference to which our photometry zeropoints are tied, the apertures for the photometry are exactly defined (although seeing obviously plays a role in smaller apertures), we illustrate how our difference images credibly remove stellar contamination, and extremely importantly, our background level is derived from a constant plus $\rho^{-3}$ fit to median ring surface brightness between radii 64–128″.
@@
Table LABEL:tab:phot gives the magnitude measured for 3I/ATLAS in four apertures drawn from stacked difference images with 2″ pixels. This flux conserving stacking was aligned on the orbit of 3I/ATLAS, without any reference to flux in the images, so the center of the central pixel corresponds to the location of 3I/ATLAS at the midpoint of the exposure. “StackID” codes site, instrument, and integer MJD night number, where a hyphen indicates the MJD range if more than $\sim$1 hour, “MJD” is the actual mean time of the exposure (1.0 July 2025 is MJD 60857.0), “N” is the number of exposures contributing to the stack, “m2” is the central 2″ pixel, “m6” is the 9 pixel $\times6$″ square aperture, “m10” is the 21 pixels of the $\times5$ pixel square less each corner pixel to make a rounded aperture of diameter $\sim$10″, “m14” is the 37 pixels of the $\times7$ pixel square removing three pixels from each corner making a rounded aperture of diameter $\sim$14″, “dm6” is the uncertainty of the “m6” magnitude (other uncertainties can be scaled by square root of area), “r” is the heliocentric distance [au], “Delta” is the Earth distance [au], “sto” is the Sun-Target-Observer angle [deg], and “HM” is the “Halley-Marcus” phase function for comets evaluated at sto, assembled by D. Schleicher [^2] https://asteroid.lowell.edu/comet/dustphase. The asterisk on 05r60868 and 05r60883 indicates potentially unreliable photometry, as described below. The typical point spread function for these observations is about 4″ full width half maximum (the $o$ and $c$ images have 1.86″/pixel and the $w$ band 1.25″/pixel) so the central pixel is significantly diminished by the seeing relative to infinite resolution. These data are available in digital form at http://astroportal.ifa.hawaii.edu/atlas/3I-photometry.